Circular No. 6408

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996ad IN ANONYMOUS GALAXY G. J. Garradd and R. H. McNaught report the discovery by R. Evans

of an apparent supernova found on a U.K. Schmidt film taken on May 9.56 UT by M. Drinkwater. McNaught notes that the new star is close to the galaxy’s center (2′′ east, 1′′ north) and difficult to see; he measured SN 1996ad (mag 18.5) on the discovery film: α = 15^{h}59^{m}05^{s}.91, δ = +19^{o}48′08′′.1 (equinox 2000.0). L. Germany (Mount Stromlo and Siding Spring Observatories) reports that the new star is present at V ∼ 18.5 on images taken with the 1-m telescope on May 17.4. A U.K. Schmidt film taken on May 25.6 by D. M. James and M. Hartley shows that SN 1996ad had faded ∼ 1 mag or more and is present only as an extension of the nucleus to the northeast. SN 1996ad is not present on earlier U.K. Schmidt films taken on 1995 July 18 and Aug. 29. McNaught further notes that the host galaxy normally shows a near-stellar nucleus, whereas at discovery there was clearly an additional stellar image adjacent to the nucleus. A foreground star (mag R = 19.0) is located ∼ 15′′ east and 5′′ north of the galaxy’s center.

COMET C/1995 O1 (HALE-BOPP) L. M. Woodney, University of Maryland (UM); M. Womack and D.

Suswal, Pennsylvania State University; J. McMullin, Submillimeter Telescope Observatory; M. F. A’Hearn, UM; M. Festou, Observatoire Midi Pyrenees; and S. A. Stern, Southwest Research Institute, report observations of H_{2}S 1(10)-1(01) and CO J(2-1) from this comet: “H_{2}S was detected for the first time in C/1995 O1 on May 21 UT at the NRAO 12-m telescope on Kitt Peak. The line was 0.6 km/s wide, had an integrated main beam brightness temperature of 0.10 K km/s, and was blueshifted by 0.15 km/s. Assuming a rotational temperature of 30 K, we derive a production rate of Q(H_{2}S) = 3.0 × 10^{26} s^{−1}. This detection was confirmed on May 23, when the line was observed to be 0.7 km/s wide and have an integrated main-beam brightness temperature of 0.17 K km/s. The CO line observed on the same day yields a production rate of Q(CO) = 4.0 × 10^{28} s^{−1}, which implies Q(H_{2}S)/Q(CO) = 0.008.”

Visual m_{1} estimates: May 18.54 UT, 7.2 (S. J. O’Meara, Volcano, HI, naked eye); 21.04, 6.8 (M. V. Zanotta, Mount Generoso, Italy, 7×42 binoculars); 24.78, 6.5 (T. Lovejoy, Jimboomba, Qld., Australia, naked eye); 26.10, 6.6 (J. Carvajal, Teruel, Spain, 10×70 binoculars).

1996 May 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green